OpenStudy (anonymous):

Simplify the expression. Write the result using positive exponents only. (-2x^3y^-4)(5x^-1y)

OpenStudy (anonymous):

(-2x^3y^-4)(5x^-1y) =(-2)(5)x^(3-1) y^(-4+1) =-10x^2 y^(-3) =-10x^2 / (y^3)